OpenStudy (anonymous):

Show all steps used to solve the problem. 5n + 6 = -4

OpenStudy (saifoo.khan):

\[5n = -4 - 6\]\[5n = -10\]

OpenStudy (saifoo.khan):

\[n = -\frac{10}{5}\]\[n = -2\]

OpenStudy (anonymous):

in the first step do you subtract 4 from botn sides i don't get it